OpenStudy (nkhanfar):

I know there is y=ax^2+bx+c, but i never saw someting like ax^2+c. How do I solve that?

OpenStudy (sunnnystrong):

You can just isolate x^2 and take the square root... so set it equal to zero... 0=ax^2+c -c=ax^2 -c/a=x^2 x= sq. rt plus/minus -c/a ... but you can also use the quadratic formula but b=0

OpenStudy (nkhanfar):

how would I apply that to y=x^2+5 @sunnnystrong?

OpenStudy (sunnnystrong):

With that quadratic I would just use the quadratic formula... because you can't take the square root of a negative number.

OpenStudy (sunnnystrong):

sorry... -b plus minus (sq. root of b^2+4ac) all divided by 2a

OpenStudy (nkhanfar):

thanks!